OpenStudy (anonymous):

Check my solution please? ^^; What are the x-intercept(s) of the graph of the equation y - x2 = -6x - 7? is it (-1, 0) and (7, 0)?

OpenStudy (anonymous):

0 - x2 = -6x -7 0 = x2 -6x - 7 (x - 7 (x + 1 x = 7 x = -1 yes ur answer is right - good job

OpenStudy (anonymous):

:D thank you!